Please help with this question. 12. (15 points) Let X be a continuous random variable with cumulative distribution function 0. F(x) = Inc. <a a<x<b bcx 1. (a
Answer #1

(a)

For continuous distribution function,

F(a) = 0

=> ln(a) = 0

=> a = 1

and , F(b) = 1

ln(b) = 1

b = e where e is exp(1)

(b)

P(X > 2) = 1 - P(X \le 2)

= 1 - F(2)

= 1 - ln(2) (as, 1 < 2 < e)

= 1 - 0.6931472

= 0.3068528

(c)

The probability density function of X is,

f(x) = F'(x) = 1/x for 1 < x < e

else f(x) = 0

(d)

E(X) = \int_{1}^{e} x f(x)~dx = \int_{1}^{e} x . (1/x)~dx = \int_{1}^{e} 1~dx = \left [ x \right ]_{1}^{e} = e-1

= 2.718282 - 1

= 1.718282